OpenStudy (anonymous):

Please help with these two problems! I'm really stuck! Which of the following is a solution to √x+10 +3=8? (The radical ends at 10) my options are x= 15 x=5 x=-5 x=-15 Identify all of the following solutions of √x+14+2=x.(the radical ends at 14) my options are x=-6 x=1 x=-6 and x=1 none of the above. Thanks in advance I would really appreciate it!

OpenStudy (saifoo.khan):

√(x+10) +3=8 √(x+10) = 5 take square on both sides. x+10 = 25

OpenStudy (anonymous):

Thank you so much !!

OpenStudy (saifoo.khan):

For the second one: √(x+14)+2=x √(x+14) = x-2 take square on both sides, x+14 = (x-2)^2 After you get 2 solutions, insert back into the equation and see if the solutions are real or not.